# CodeAssistBench (CAB): Dataset & Benchmarking for Multi-turn Chat-Based Code Assistance

### v4

@@ Abstract
Programming assistants powered by large language models have improved dramatically, yet existing benchmarks still evaluate them in narrow code-generation settings. Recent efforts such as InfiBench and StackEval rely on Stack Overflow questions and remain limited to single-turn interactions, manually curated data, and isolated snippets rather than full project environments. We introduce CodeAssistBench (CAB), the first benchmark for evaluating multi-turn, project-grounded programming assistance at scale. CAB automatically constructs datasets from GitHub issues tagged as questions, using an LLM-driven pipeline that filters noise, extracts runnable contexts, builds executable containers, and verifies environment correctness. This enables continuous, automated expansion across diverse repositories without manual intervention. Using CAB, we create a testbed of 3,286 real-world issues across 214 repositories, spanning seven languages. Evaluating state-of-the-art models reveals a substantial gap: while models achieve 70–83% accuracy on Stack Overflow–style questions, they solve only 16.49% of CAB issues from post-training-cutoff repositories. On a manually validated subset of 149 issues, top models such as Claude Sonnet 4.5 reach only 12.08% correctness. These results highlight a fundamental challenge: current LLMs struggle to provide assistance in realistic, project-specific contexts despite strong performance on traditional Q&A benchmarks. CAB provides a scalable, reproducible framework for advancing research in multi-turn, codebase-grounded programming agents. The benchmark and pipeline are fully automated and publicly available at https://github.com/amazon-science/CodeAssistBench/.
@@ 1 Introduction
Large language models (LLMs) are increasingly integrated into modern programming workflows, supporting tasks ranging from code generation to debugging and repository navigation. While benchmarks have progressed from isolated synthesis tasks (HumanEval chen2021evaluating, MBPP austin2021program) to repository-level maintenance settings (SWE-Bench jimenez2023swe, BigCodeBench zhuo2024bigcodebench), they still capture only a narrow slice of how developers seek assistance in practice. Recent multi-turn benchmarks such as ConvCodeWorld han2025convcodeworld, MINT wang2023mint, and TICODER fakhoury2024ticoder move toward conversational evaluation, yet they primarily emphasize code synthesis and assume stable, well-defined contexts.
However, real-world programming assistance extends far beyond writing code. A 2024 Stack Overflow survey of 34,168 developers reports that while 75.7% seek AI for code generation, even higher percentages desire support for searching for answers (77.9%), debugging and troubleshooting (77.3%), and understanding unfamiliar codebases (73.6%) stackoverflow2024aisurvey. These tasks require iterative clarification, environment-aware reasoning, and integration with project-specific details—capabilities that current benchmarks largely overlook.
Recent efforts such as InfiBench li2024infibench and StackEval shah2024stackeval incorporate Stack Overflow data to better reflect developer Q&A scenarios, but they remain limited to single-turn interactions over isolated snippets and require substantial manual curation. Critically, they do not model how developers and assistants interact over multiple turns: trying suggestions, encountering new errors, revising assumptions, and progressively forming shared understanding.
@@
Our evaluation of leading LLMs reveals a striking capability gap. While recent models achieve 70–83% accuracy on Stack Overflow-style questions, they solve only 16.49% of CAB issues from post-training-cutoff repositories. Even on a manually validated subset of 149 issues, top models reach only 12.08% correctness. These results highlight a fundamental challenge: despite impressive performance on traditional Q&A benchmarks, LLMs struggle to provide reliable assistance in realistic, multi-turn, project-specific contexts.
@@ 2 Related Work
Code Generation Benchmarks. Traditional evaluations of programming assistants focus on generating functionally correct code from well-specified prompts. HumanEval chen2021evaluating, MBPP austin2021program, and CodeContests li2022competition assess single-shot code synthesis, while more advanced benchmarks such as SWE-Bench jimenez2023swe and SWE-PolyBench rashid2025swepolybenchmultilanguagebenchmarkrepository use real GitHub issues to evaluate repository-level maintenance tasks. These benchmarks provide robust execution-based correctness metrics but remain predominantly single-turn and do not capture the iterative, conversational nature of real developer interactions.
Multi-turn Programming and Conversational Benchmarks. Several benchmarks incorporate iterative interactions. ConvCodeWorld han2025convcodeworld, MINT wang2023mint, and TICODER fakhoury2024ticoder simulate developer–assistant dialogues but typically assume stable, fully specified environments and remain centered on code synthesis tasks. More general conversational evaluations—such as MT-Bench zheng2023judging, AlpacaEval dubois2023alpacafarm, and BotChat duan-etal-2024-botchat—measure dialogue quality but are not targeted at technical problem-solving in programming contexts. In contrast, we evaluate multi-turn, task-grounded interactions where assistance depends on project state, ambiguous context, and evolving user needs.
Programming Q&A Benchmarks. StackEval shah2024stackeval and InfiBench li2024infibench use Stack Overflow data to better reflect practical developer questions and emphasize explanatory reasoning. However, they remain limited to single-turn settings, require substantial manual curation, and lack integration with executable project environments. These constraints prevent them from capturing the iterative reasoning, context reconstruction, and environment-aware troubleshooting that dominate real-world programming assistance.
@@ 3.1.1 Repository Collection
To prioritize repositories with active developer support communities, we define a community score $\mathrm{CS}(r)=Q(r)+H(r)$, where $Q(r)$ and $H(r)$ represents the number of closed issues labeled with “question” and “help wanted” respectively. We then select the top-$K$ repositories ranked by $\mathrm{CS}(r)$ (breaking ties by star count) to form our repository set $R_{K}$, where $K$ is a user-defined parameter.
@@ 3.1.2 Issue Filtering
For each repository $r\in R_{K}$, let $I_{r}$ be the set of issues associated with it. We fetch only successfully closed issues using the GitHub Issues API. Each issue is represented as:
$i=(\mathrm{title}(i),\,\mathrm{body}(i),\,\mathrm{messages}(i)),\quad\mathrm{messages}(i)=\{m_{i,1},m_{i,2},\dots\}$ (2)
@@ 3.1.3 Data Preparation
Build Environment Generation. For issues requiring environment-specific testing, we automatically generate Docker configurations by analyzing repository artifacts using Sonnet 3.7 claude37sonnet with a structured prompt (Appendix A.3) that identifies the commit $\mathit{sha}_{i}$ closest to the issue creation timestamp, clones the repository at that commit, and extracts key artifacts (README content, Dockerfiles, GitHub workflows, file structure). It then generates and tests candidate build scripts until finding a successful configuration $e_{i}$.
@@ 3.2 Evaluation Framework and Implementation
Figure 3: CAB evaluation pipeline. A simulated user chats with the Maintainer Agent, which can run code in an optional build sandbox; the interaction continues until the user is satisfied or reaches the maximum turn limit. Once the dialogue ends, an LLM-judge grades the exchange against satisfaction conditions extracted from the original GitHub issue. This pipeline enables realistic assessment of programming assistance in context-rich, project-specific scenarios.
CAB’s evaluation framework simulates realistic programming assistance interactions through a multi-agent system with three distinct roles: a User, a Maintainer, and a Judge (see Appendix Figure 3).
@@
Maintainer Agent. The maintainer agent represents the LLM being evaluated. Given the user’s question and contextual information, the model must analyze the problem within the provided codebase context, execute commands in the containerized environment when necessary, generate helpful responses, and adapt its approach based on user feedback. To balance exploration with efficiency, the maintainer can take up to 5 exploration steps per response (configurable), where each step may involve file operations or command execution.
@@ 4.1 Statistics of Dataset Generation Pipeline
Table 1: Summary of issue filtering, Docker requirement detection, build outcomes, and final issue retention across programming languages and repository cohorts in our dataset generation pipeline.
@@
We applied our automated pipeline (Section 3.1) to two repository cohorts: an All-Time set of 700 top-starred repositories (100 per language) without creation-date limits, and a Recent set of 3,500 repositories (500 per language) created after November 2024. After applying license and community filters, we retain 770 repositories for downstream processing. These produced 25,656 raw GitHub issues tagged with question or help-wanted. We apply a layered filtering process, regex heuristics followed by an LLM classifier, to remove low-signal or ambiguous issues.
Across both cohorts, we retained 3,286 multi-turn issues from 214 repositories, including 238 fully validated Docker environments. Notably, Recent repositories achieved a 97.6% build success rate compared to 78.2% for All-Time repositories (details in Appendix F). In total, our pipeline executed 44,628 Sonnet 3.7 calls end-to-end without any manual intervention. Issue-length distributions (Figure 5) show that over half of all issues involve multi-turn threads, particularly in Python, C++, and TypeScript—highlighting the need for multi-step reasoning benchmarks such as CAB.
@@ 4.2.4 Qualitative Example
Figure 4 illustrates a GitHub issue where the user asks how to remap ports for both the main application and the Cloudflare proxy to avoid conflicts with other apps. It presents two contrasting model stories: Haiku 3.5 provides an incomplete response, either omitting clear proxy-port guidance or introducing incorrect configurations, while ChatGPT 4.1 Mini offers a successful solution - clearly explaining how to change the main app’s port, noting that the proxy port doesn’t need to be exposed, and providing a conflict-free docker-compose.yml. A simulated user agent drives each LLM through a multi-turn dialogue, asking follow-up questions until all satisfaction conditions are met or a predefined turn limit is reached. Where applicable, the agent also validates the recommended configuration using recorded build outcomes. Once the conversation concludes, an automated LLM-Judge assigns a final label - Correct, Partially Correct, or Incorrect - based on comparisons with the original GitHub thread, execution results, and the predefined satisfaction conditions.
@@ 4.2.6 Verbosity Analysis
Beyond correctness, we assess whether models provide appropriately concise explanations. Figure 7 (Appendix) summarizes verbosity classifications across all models. Overall, models tend to err on the side of over-explanation: verbose responses account for 40–60% of outputs, while terse responses are rare (<5%). Sonnet 3.7 (Think) shows the most balanced behavior, producing the highest proportion of appropriately concise answers, while Haiku 3.5 and Llama 3.3 70B are the most consistently verbose. Interestingly, verbosity increases on recent repositories, likely due to higher uncertainty and evolving APIs. These results suggest that current LLMs prioritize thoroughness over brevity, and that verbosity patterns meaningfully vary across model families.
@@ 4.3.1 Judge Validation
To evaluate judge robustness across model upgrades, we additionally ran three successive versions of Anthropic’s Sonnet family—Sonnet 3.7, Sonnet 4.0, and Sonnet 4.5—as maintainer agents on the 149 issues where human annotators reached full agreement. Individually, Sonnet 3.7 produced 12 correct, 45 partially correct, and 87 incorrect responses; Sonnet 4.0 produced 18 correct, 38 partially correct, and 92 incorrect responses; and Sonnet 4.5 produced 18 correct, 48 partially correct, and 82 incorrect responses. The consistency of these distributions across model generations indicates that judge behavior is stable with respect to model choice, and that the verified subset is representative for downstream analyses.
Comparing the automated judge to human annotators, the LLM judge achieved 65.92% average agreement—84.2% of the inter-human baseline. Agreement was highest on objective dimensions such as verbosity (93.2% of human-level agreement) and lower on subjective dimensions such as technical correctness (86.7% of human-level agreement). Based on these results, we release the 149-issue verified subset (84.7% of the validation set) as a high-quality set of gold-standard labels for future research. Full methodology, annotation protocols, and judge–human comparison details are provided in Appendix B.
@@ 4.3.2 Error Category Analysis
To better characterize how models fail across different problem types, we categorized the 310 cases from our human annotation study (Section B) into seven coarse error types using LLM-assisted classification with Sonnet 4.5. Logic errors (28.8%) and environment configuration issues (28.5%) dominate the dataset, while performance-related issues exhibit the highest annotator–judge agreement (75.0%). The full error taxonomy, per-model breakdowns, and implications for future benchmark design are provided in Appendix K.
@@ 5 Limitations
While CodeAssistBench (CAB) offers a realistic and scalable framework for evaluating multi-turn programming assistance, it has several limitations. These include conservative condition extraction, limited evaluation coverage, templated user behavior, language scope constraints, and lack of statistical testing. We provide a full discussion in Appendix C.
@@ 6 Conclusion and Future Work
Our study introduces CodeAssistBench (CAB), a fully automated benchmark for evaluating multi-turn programming assistance grounded in real-world developer interactions. Unlike prior benchmarks that emphasize single-turn or synthetic tasks, CAB simulates realistic conversations over full codebases using containerized environments and issue-specific satisfaction criteria. Through extensive experiments, we find that state-of-the-art language models struggle with complex, multi-turn dialogues - particularly on recent repositories, where even the best models achieve only 12.08% correctness on the verified subset. CAB’s automated pipeline enables scalable dataset expansion while supporting rigorous, reproducible evaluation across languages, project types, and time periods, offering a more faithful measure of model capabilities than existing benchmarks.
@@ 3. APPROPRIATE VERBOSITY
Listing 5: Prompt for LLM-Judge
@@ B.1 Validation Methodology
We then generated AI responses using two state-of-the-art models for these 176 issues: GPT-4.1-mini and Claude Sonnet 3.5. Initially, both models generated responses for most issues (134 issues received responses from both models to enable comparative analysis). However, not all AI-generated responses were suitable for human annotation. We excluded responses where the AI output contained unicode character encoding issues or null fields in critical parts, which would have prevented annotators from accurately evaluating the responses. Specifically, 40 GPT responses and 38 Sonnet responses were excluded due to these data quality issues.
After this filtering, we retained 310 usable AI responses for human evaluation: 154 from GPT-4.1-mini and 156 from Claude Sonnet 3.5. Each of these 310 responses was independently evaluated by two expert annotators, yielding 620 total annotation records. The final language-wise distribution of validated AI responses per model is shown in Table 3.
Table 3: Language distribution of validated AI responses
Language GPT-4.1-mini Claude Sonnet 3.5
@@
Two senior software engineers (Human1 and Human2) with 3+ years of experience each independently evaluated all 310 AI responses. Critically, annotators used the exact same evaluation prompts and criteria as the LLM judges to ensure fair comparison and minimize methodological confounds. Each response was evaluated across three dimensions:
@@
Annotators worked independently without communication, and each AI response was evaluated by both annotators, yielding 620 annotation records (310 responses $\times$ 2 annotators).

### v5

@@ Abstract
Programming assistants powered by large language models have improved dramatically, yet existing benchmarks still evaluate them in narrow code-generation settings. Recent efforts such as InfiBench and StackEval rely on Stack Overflow questions and remain limited to single-turn interactions, manually curated data, and isolated snippets rather than full project environments. We introduce CodeAssistBench (CAB), the first benchmark for evaluating multi-turn, project-grounded programming assistance at scale. CAB automatically constructs datasets from GitHub issues tagged as questions, using an LLM-driven pipeline that filters noise, extracts runnable contexts, builds executable containers, and verifies environment correctness. This enables continuous, automated expansion across diverse repositories without manual intervention. Using CAB, we create a testbed of 3,286 real-world issues across 214 repositories, spanning seven languages. Evaluating state-of-the-art models reveals a substantial gap: while models achieve 70–83% accuracy on Stack Overflow–style questions, they solve only 7.22–16.49% of CAB issues from post-training-cutoff repositories. These results highlight a fundamental challenge: current LLMs struggle to provide assistance in realistic, project-specific contexts despite strong performance on traditional Q&A benchmarks. CAB provides a scalable, reproducible framework for advancing research in multi-turn, codebase-grounded programming agents. The benchmark and pipeline are fully automated and publicly available at https://github.com/amazon-science/CodeAssistBench/.
@@ 1 Introduction
Large language models (LLMs) are increasingly integrated into modern programming workflows, supporting tasks ranging from code generation to debugging and repository navigation. While benchmarks have progressed from isolated synthesis tasks (HumanEval Chen et al. (2021), MBPP Austin et al. (2021)) to repository-level maintenance settings (SWE-Bench Jimenez et al. (2023), BigCodeBench Zhuo et al. (2024)), they still capture only a narrow slice of how developers seek assistance in practice. Recent multi-turn benchmarks such as ConvCodeWorld Han et al. (2025), MINT Wang et al. (2023), and TICODER Fakhoury et al. (2024) move toward conversational evaluation, yet they primarily emphasize code synthesis and assume stable, well-defined contexts.
However, real-world programming assistance extends far beyond writing code. A 2024 Stack Overflow survey of 34,168 developers reports that while 75.7% seek AI for code generation, even higher percentages desire support for searching for answers (77.9%), debugging and troubleshooting (77.3%), and understanding unfamiliar codebases (73.6%) Stack Overflow (2024). These tasks require iterative clarification, environment-aware reasoning, and integration with project-specific details—capabilities that current benchmarks largely overlook.
Recent efforts such as InfiBench Li et al. (2024) and StackEval Shah et al. (2024) incorporate Stack Overflow data to better reflect developer Q&A scenarios, but they remain limited to single-turn interactions over isolated snippets and require substantial manual curation.
@@
Our evaluation of leading LLMs reveals a striking capability gap. While recent models achieve 70–83% accuracy on Stack Overflow-style questions, they solve only 16.49% of CAB issues from post-training-cutoff repositories. These results highlight a fundamental challenge: despite impressive performance on traditional Q&A benchmarks, LLMs struggle to provide reliable assistance in realistic, multi-turn, project-specific contexts.
@@ 2 Related Work
Code Generation Benchmarks. Traditional evaluations of programming assistants focus on generating functionally correct code from well-specified prompts. HumanEval Chen et al. (2021), MBPP Austin et al. (2021), and CodeContests Li et al. (2022) assess single-shot code synthesis, while more advanced benchmarks such as SWE-Bench Jimenez et al. (2023) and SWE-PolyBench Rashid et al. (2025) use real GitHub issues to evaluate repository-level maintenance tasks. These benchmarks provide robust execution-based correctness metrics but remain predominantly single-turn and do not capture the iterative, conversational nature of real developer interactions.
Multi-turn Programming and Conversational Benchmarks. Several benchmarks incorporate iterative interactions. ConvCodeWorld Han et al. (2025), MINT Wang et al. (2023), and TICODER Fakhoury et al. (2024) simulate developer–assistant dialogues but typically assume stable, fully specified environments and remain centered on code synthesis tasks. More general conversational evaluations—such as MT-Bench Zheng et al. (2023), AlpacaEval Dubois et al. (2023), and BotChat Duan et al. (2024)—measure dialogue quality but are not targeted at technical problem-solving in programming contexts. In contrast, we evaluate multi-turn, task-grounded interactions where assistance depends on project state, ambiguous context, and evolving user needs.
Programming Q&A Benchmarks. StackEval Shah et al. (2024) and InfiBench Li et al. (2024) use Stack Overflow data to better reflect practical developer questions and emphasize explanatory reasoning. However, they remain limited to single-turn settings, require substantial manual curation, and lack integration with executable project environments. These constraints prevent them from capturing the iterative reasoning, context reconstruction, and environment-aware troubleshooting that dominate real-world programming assistance.
@@ 3.1.1 Repository Collection
To prioritize repositories with active developer support communities, we define a community score $\mathrm{CS}(r)=Q(r)+H(r)$, where $Q(r)$ and $H(r)$ represents the number of closed issues labeled with “question” and “help wanted” respectively. We then select the top-$N$ repositories ranked by $\mathrm{CS}(r)$ (breaking ties by star count) to form our repository set $R_{N}$, where $N$ is a user-defined parameter.
@@ 3.1.2 Issue Filtering
For each repository $r\in R_{N}$, let $I_{r}$ be the set of issues associated with it. We fetch only successfully closed issues using the GitHub Issues API. Each issue is represented as:
$i=(\mathrm{title}(i),\,\mathrm{body}(i),\,\mathrm{messages}(i)),\quad\mathrm{messages}(i)=(m_{i,1},m_{i,2},\dots)$ (2)
@@
To construct turns, we group consecutive messages from the same role (author or maintainer) into a single logical message, then pair each author segment with the subsequent maintainer segment. Unpaired trailing messages (e.g., a final “thanks” from the author) are discarded.
@@ 3.1.3 Data Preparation
Build Environment Generation. For issues requiring environment-specific testing, we automatically generate Docker configurations by analyzing repository artifacts using Sonnet 3.7 Anthropic (2025) with a structured prompt (Appendix A.3) that identifies the commit $\mathit{sha}_{i}$ closest to the issue creation timestamp, clones the repository at that commit, and extracts key artifacts (README content, Dockerfiles, GitHub workflows, file structure). It then generates and tests candidate build scripts until finding a successful configuration $e_{i}$.
@@ 3.2 Evaluation Framework and Implementation
Figure 3: CAB evaluation pipeline. A simulated user chats with the Maintainer Agent, which can run code in an optional build sandbox; the interaction continues until the user is satisfied or reaches the maximum turn limit. Once the dialogue ends, an LLM judge grades the exchange against satisfaction conditions extracted from the original GitHub issue. This pipeline enables realistic assessment of programming assistance in context-rich, project-specific scenarios.
CAB’s evaluation framework simulates realistic programming assistance interactions through a multi-agent system with three distinct roles: a User, a Maintainer, and a Judge (see Figure 3).
@@
Maintainer Agent. The maintainer agent represents the LLM being evaluated. Given the user’s question and contextual information, the model must analyze the problem within the provided codebase context, execute commands in the containerized environment when necessary, generate helpful responses, and adapt its approach based on user feedback.
@@ 4.1 Statistics of Dataset Generation Pipeline
Table 1: Summary of issue filtering, Docker requirement detection, build outcomes, and final issue retention across programming languages and repository cohorts in our dataset generation pipeline. The 56-issue difference between filtered issues (3,342) and final retained issues (3,286) corresponds to Docker-required issues where environment builds failed (294 Docker-required - 238 successful builds == 56 excluded).
@@
We applied our automated pipeline (Section 3.1) to two repository cohorts: an All-Time set of 700 top-starred repositories (100 per language) without creation-date limits, and a Recent set of 3,500 repositories (500 per language) created after November 2024. After applying license and community filters, we retain 770 repositories for downstream processing. These produced 25,656 raw GitHub issues tagged with question or help-wanted. We apply a layered filtering process, regex heuristics followed by an LLM classifier, to remove low-signal or ambiguous issues. Issues requiring Docker environments are excluded if automated build validation fails.
Across both cohorts, we retained 3,286 multi-turn issues from 214 repositories, including 238 fully validated Docker environments. Many repositories yielded no qualifying multi-turn issues after filtering, resulting in 214 repositories contributing at least one benchmark instance. Notably, Recent repositories achieved a 97.6% build success rate compared to 78.2% for All-Time repositories (details in Appendix F). In total, our pipeline executed 44,628 Sonnet 3.7 calls end-to-end without any manual intervention. Issue-length distributions (Figure 5) show that over half of all issues involve multi-turn threads, particularly in Python, C++, and TypeScript—highlighting the need for multi-step reasoning benchmarks such as CAB.
@@ 4.2.4 Qualitative Example
Figure 4 illustrates a GitHub issue where the user asks how to remap ports for both the main application and the Cloudflare proxy to avoid conflicts with other apps. It presents two contrasting model stories: Haiku 3.5 provides an incomplete response, either omitting clear proxy-port guidance or introducing incorrect configurations, while ChatGPT 4.1 Mini offers a successful solution - clearly explaining how to change the main app’s port, noting that the proxy port doesn’t need to be exposed, and providing a conflict-free docker-compose.yml. A simulated user agent drives each LLM through a multi-turn dialogue, asking follow-up questions until all satisfaction conditions are met or a predefined turn limit is reached. Where applicable, the agent also validates the recommended configuration using recorded build outcomes. Once the conversation concludes, an automated LLM judge assigns a final label - Correct, Partially Correct, or Incorrect - based on comparisons with the original GitHub thread, execution results, and the predefined satisfaction conditions.
@@ 4.2.6 Verbosity Analysis
Figure 7 (Appendix) summarizes verbosity classifications. Models tend toward over-explanation: verbose responses account for 40–60% of outputs. Sonnet 3.7 (Think) shows the most balanced behavior, while Haiku 3.5 and Llama 3.3 70B are most verbose. Verbosity increases on recent repositories, likely due to higher uncertainty.
@@ 4.3.1 Judge Validation
Comparing the automated judge to human annotators, the LLM judge achieved 65.92% average agreement—84.2% of the inter-human baseline. Agreement was highest on objective dimensions such as verbosity (93.2% of human-level agreement) and lower on subjective dimensions such as technical correctness (86.7% of human-level agreement). Full methodology, annotation protocols, and judge–human comparison details are provided in Appendix B.
@@ 4.3.2 Error Category Analysis
To better characterize how models fail across different problem types, we categorized the 310 cases from our human annotation study (Section B) into seven coarse error types using LLM-assisted classification with Sonnet 3.7. Logic errors (28.8%) and environment configuration issues (28.5%) dominate the dataset, while performance-related issues exhibit the highest annotator–judge agreement (75.0%). The full error taxonomy, per-model breakdowns, and implications for future benchmark design are provided in Appendix K.
@@ 5 Limitations
While CodeAssistBench (CAB) offers a realistic and scalable framework for evaluating multi-turn programming assistance, several limitations remain, including conservative condition extraction, limited evaluation coverage, templated user behavior, and language scope constraints. Additionally, our LLM-based judge achieves 65.92% agreement with human annotators (84.2% of the 78.28% inter-human baseline), suggesting multi-turn evaluation introduces alignment challenges compared to single-turn settings; see Appendix C for full discussion.
@@ 6 Conclusion and Future Work
Our study introduces CodeAssistBench (CAB), a fully automated benchmark for evaluating multi-turn programming assistance grounded in real-world developer interactions. Unlike prior benchmarks that emphasize single-turn or synthetic tasks, CAB simulates realistic conversations over full codebases using containerized environments and issue-specific satisfaction criteria. Through extensive experiments, we find that state-of-the-art language models struggle with complex, multi-turn dialogues - particularly on recent repositories. CAB’s automated pipeline enables scalable dataset expansion while supporting rigorous, reproducible evaluation across languages, project types, and time periods, offering a more faithful measure of model capabilities than existing benchmarks.
@@ 3. APPROPRIATE VERBOSITY
Listing 5: Prompt for LLM judge
@@ B.1 Validation Methodology
We then generated agent responses using two state-of-the-art models for these 176 issues: ChatGPT 4.1 Mini and Claude Sonnet 3.7. Initially, both models generated responses for most issues (134 issues received responses from both models to enable comparative analysis). However, not all AI-generated responses were suitable for human annotation. We excluded responses where the agent output contained unicode character encoding issues or null fields in critical parts, which would have prevented annotators from accurately evaluating the responses. Specifically, 40 GPT responses and 38 Sonnet responses were excluded due to these data quality issues.
After this filtering, we retained 310 usable agent responses for human evaluation: 154 from ChatGPT 4.1 Mini and 156 from Claude Sonnet 3.7. Each of these 310 responses was independently evaluated by two expert annotators, yielding 620 total annotation records. The final language-wise distribution of validated agent responses per model is shown in Table 3.
Table 3: Language distribution of validated agent responses
Language ChatGPT 4.1 Mini Claude Sonnet 3.7
@@
Two senior software engineers (Human1 and Human2) with 3+ years of experience each independently evaluated all 310 agent responses. Critically, annotators used the exact same evaluation prompts and criteria as the LLM judges to ensure fair comparison and minimize methodological confounds. Each response was evaluated across three dimensions:
@@
Annotators worked independently without communication, and each agent response was evaluated by both annotators, yielding 620 annotation records (310 responses $\times$ 2 annotators).